A carpenter used 54ft of molding in three pieces to trim a garage door. If the long piece was 4ft longer than three times the length of each shorter piece, then how long was each piece?
Answer by ikleyn(52781)

.
Let x be the length of each of the two shorter pieces.


Then the length of the longer piece is  3x+4, and the equation for the total is


    x + x + (3x + 4) = 54.


Simplify and solve for x.


    5x + 4 = 54

    5x = 54 - 4 = 50.

     x          = 50/5 = 10.


ANSWER.  Both short pieces are 10 ft long.  The long piece is 3*10+4 = 34 ft long.

Solved.